OpenStudy (anonymous):

find the laplace inverse transform F(s)=(s+2)/(s^2-3s+4)

OpenStudy (mathmale):

What have you done in previous problems like this one? What would the first steps be?

OpenStudy (unklerhaukus):

complete the square in the dominator

OpenStudy (mathmale):

Yes, but for what purpose? What would the very first goal be in attacking this problem, and why?

OpenStudy (unklerhaukus):

if you complete the square in the denominator you can apply the shift theorem

OpenStudy (anonymous):

To check your answer, visit http://www.wolframalpha.com/input/?i=inverse+lapalace+transform+%28s%2B2%29%2F%28s^2-3s%2B4%29

OpenStudy (anonymous):

\[inverse L=s/((s-3/2)^2+7/4) +2/((s-3/2)^2+7/4)... then ? im stuck\]

OpenStudy (mathmale):

I'd suggest you look at this table of Laplace Transforms and determine for yourself which one or two of the transform pairs would be best suited to apply to your Laplace transform problem. http://tutorial.math.lamar.edu/Classes/DE/Laplace_Table.aspx

OpenStudy (mathmale):

I would write your particular Laplace transform as \[F(s) = \frac{ s+2 }{ (s-\frac{ 3 }{ 2 })^2+(\frac{ \sqrt{7} }{ 2 } )^2}\]

OpenStudy (mathmale):

Think: How could you modify this Laplace transform so that you could use one or more of the Laplace transform pairs in the table to find the inverse Laplace transform of F(s)?

OpenStudy (mathmale):

Were it up to me, I'd use the pairs #19 and #20. Good luck!